# Stiefel-Whitney Classes for Finite Symplectic Groups

### v1

@@ Abstract.
Stiefel-Whitney classes, Symplectic Groups, Finite groups of Lie type
@@ 1. Introduction
This paper is part of a project to understand SWCs for finite groups of Lie type. Let $q$ be an odd prime power throughout. A formula determining SWCs for $G=\operatorname{GL}(n,q)$ was discovered in [GJar], for $G=\operatorname{SL}(2,q)$ in [MS23] and for $G=\operatorname{SL}(2n+1,q)$ in [MS24]. In this paper we find similar expressions when $G=\operatorname{Sp}(2n,q)$.
Write $\operatorname{H}^{*}(G)$ for the mod $2$ cohomology $\operatorname{H}^{*}(G,\mathbb{F}_{2})$, and $\operatorname{H}^{*}_{\operatorname{SW}}(G)$ for the subalgebra of $\operatorname{H}^{*}(G)$ generated by SWCs of orthogonal representations. The diagonal matrices in $G$ with eigenvalues ±$\pm 1$ form a subgroup we denote by $Z_{X}$. (It is the center of another subgroup $X$ which we will encounter later.) The mod $2$ cohomology of $Z_{X}$ is a polynomial algebra:
@@
Since no information is lost by restriction, we will express SWCs in terms of the $\mathcal{E}_{k}$. For $0\leq i\leq n$, let $g_{i}\in Z_{X}$ with $-1$ having multiplicity $2i$ as an eigenvalue, and $1$ having multiplicity $2(n-i)$. For example $g_{0}=\mathbbm{1}$, the identity matrix. Given a representation $\pi$, write $\chi_{\pi}$ for its character. Here is our “universal” formula for the 4th and 8th SWCs:
@@
$w_{8}(\pi)=r_{1}\mathcal{E}_{2}+\left(\binom{r_{1}}{2}+\binom{r_{2}}{2}\right)\mathcal{E}_{1}^{2}\quad\text{ for $n\geq 2$}.\$
@@
A genera formula for the total SWCs $w(\pi)$ is found in Theorem 3.6, although it takes some computation to extract individual SWCs $w_{k}(\pi)$ from this. For example:
@@
Let $m\geq n$. The restriction map $\iota_{n}^{*}:\operatorname{H}^{i}(G_{m})\to\operatorname{H}^{i}(G_{n-1})$ is injective for $i<4n-1$.
(Compare [Nak60, Corollary 6.7] and [GJar, Theorems 3, 8].)
@@
Acknowledgments. Part of this paper comes out of the first author’s Ph.D. thesis [Mal22] at IISER Pune, during which she was supported by a Ph.D. fellowship from the Council of Scientific and Industrial Research, India. Both authors thank the Chennai Mathematical Institute for its hospitality during many visits.
@@ 2. Notations and Preliminaries
As this paper is a continuation of [MS23], we use the same notations and conventions, which we now review.
@@ 2.1. Representations
Let $G$ be a finite group. All the representations $(\pi,V)$ considered in this paper (before the Appendix) are complex finite dimensional. Let $\operatorname{Irr}(G)$ be the set of isomorphism classes of irreducible representations of $G$. Write $(\pi^{\vee},V^{\vee})$ for the dual representation. If $H$ is a subgroup of $G$, write $\operatorname{res}^{G}_{H}\pi$ or $\pi|_{H}$ for the restriction of $\pi$ to $H$. A linear character $\chi$ of $G$ is a degree 1 representation. We call $\chi$ quadratic provided $\chi^{2}=1$.
We say $\pi$ is orthogonal (resp., symplectic), provided there exists a non-degenerate $G$-invariant symmetric (resp., antisymmetric) bilinear form $B:V\times V\to\mathbb{C}$. When $\pi$ is self-dual and irreducible, it is either orthogonal or symplectic. In this case, the Frobenius-Schur Indicator $\varepsilon(\pi)$ is a sign defined as $1$ when $\pi$ is orthogonal, and $-1$ when $\pi$ is symplectic. Whereas it is 00, when $\pi$ is not self-dual.
One can symmetrize a general $(\pi,V)$ by defining $S(\pi):=\pi\oplus\pi^{\vee}$ on the vector space $V\oplus V^{\vee}$. Under the symmetric $G$-invariant bilinear map $B$ on $V\oplus V^{\vee}$ as $B((v,\alpha),(w,\beta))=\langle\alpha,w\rangle+\langle\beta,v\rangle$, the representation $S(\pi)$ is orthogonal. We call $S(\pi)$ the symmetrization of $\pi$. (In [Sch12, Definition 4.3 ] it is called the hyperbolic space on $V$.)
@@ 2.3. Characteristic Classes
known as the $i$th Stiefel-Whitney Class (SWC) of $\pi$. Their sum $w(\pi)=w_{0}(\pi)+w_{1}(\pi)+\ldots$ is called the total SWC of $\pi$. We refer the reader to [MS23, Section 2.3] for detailed description.
Also, associated to a complex representation $\pi$ of $G$ are cohomology classes $c_{i}(\pi)\in\operatorname{H}^{2i}(G,\mathbb{Z}),$ called Chern classes (CCs). Their sum $c(\pi)=c_{0}(\pi)+c_{1}(\pi)+c_{2}(\pi)+\ldots\in\operatorname{H}^{*}(G,\mathbb{Z})$ is called the total Chern class of $\pi$. We have
@@
where $\kappa:\operatorname{H}^{*}(G,\mathbb{Z})\rightarrow\operatorname{H}^{*}(G,\mathbb{F}_{2})$ is the coefficient homomorphism of cohomology. (See [GJ22, Lemma 1], based on [MS16, Problem 14-B] for proof.) So we interpret $w(S(\pi))$ as the “mod $2$ Chern class” of $\pi$.
@@
Proposition 2.1 ([May99], Chapter 23, Section 6).
@@ 2.4. Symmetric Functions
Whereas for $k>n$, these are defined to be 00. For example, $\mathcal{E}_{1}(x_{1},x_{2})=x_{1}+x_{2}$ and $\mathcal{E}_{2}(x_{1},x_{2})=x_{1}x_{2}$. Of course,
@@ 2.5. Cyclic Groups
It is known [KT67] that for $k\equiv 2\pmod{4}$, $\operatorname{H}^{*}(C_{k})=\mathbb{F}_{2}[v]$ where $v=w_{1}(\operatorname{sgn})$. In this case, (see [MS23, Lemma 2.5] for instance) for a representation $\pi$,
@@ 2.6. OIRs of a Direct Product of Groups
for $(g_{1},\ldots,g_{n})\in G$ and $v_{1}\otimes\cdots\otimes v_{n}\in V_{1}\otimes\cdots\otimes V_{n}.$ (See [FH13, Exercise 2.36, page 24].) If each $\pi_{i}$ is irreducible, then so is $\pi$. All irreducible representations of $G$ decompose in this manner. The Frobenius-Schur indicator of $\pi$ is given by
@@ 2.7. Quaternion Group
Let $Z$ be the center of $Q$, which is {±$\{\pm 1\}$. For a linear character $\theta$ of $Z^{n}$, write $[\theta,\pi]$ for the multiplicity of $\theta$ in $\pi|_{Z^{n}}$.
@@
Let $\pi$ be an orthogonal representation of $Q^{n}$. Let $\theta$ be a non-trivial linear character of $Z^{n}$. Then $[\theta,\pi]$ is a multiple of $4$.
@@
$\theta_{i}=\begin{cases}\operatorname{sgn},&i\in I_{\pi}\\ 1&i\notin I_{\pi}.\\ \end{cases}$
Then $\pi|_{Z^{n}}\cong 2^{r}\theta_{\pi}$, since $\rho|_{Z}=\operatorname{sgn}\oplus\operatorname{sgn}$. If $\pi$ is irreducible orthogonal, then $r$ is even by Lemma 2.2 and therefore $4$ divides $[\theta_{\pi},\pi]$, and other $[\theta,\pi]=0$.
@@ 2.8. $S_{n}$-invariant Representations of Elementary Abelian $2$-groups
is $S_{n}$-invariant, of degree $\binom{n}{k}$. For a polynomial $f$, write $[f]_{i}$ for the coefficient of the degree $i$ term of $f$; in other words so that $f(x)=\sum_{i}[f]_{i}x^{i}$.
Lemma 2.4 ([GJar], Proposition 2).
@@
where $m_{k}(\sigma)$ are certain non-negative integers. The coefficients $m_{k}(\sigma)$ can be expressed in terms of character values of $\sigma$ at the elements of $E$. For $0\leq k\leq n$, put $\vartheta_{k}=e_{1}+\ldots+e_{k}$. (In particular, $\vartheta_{0}=0$.) Then,
@@
From Lemma 2.4, the character value $\chi_{\sigma_{i}}(\vartheta_{k})$ is the coefficient of $x^{i}$ in the expression $(1-x)^{k}(1+x)^{n-k}$. Following [GJar, Theorem 7] we have
@@
$m_{k}(\sigma)=\begin{cases}0&\text{ when }k\text{ is odd}\\[5.0pt] \cfrac{1}{2^{n-1}}\sum\limits_{i=0}^{\frac{n-1}{2}}\chi_{\sigma_{i}}(\vartheta_{k})\chi_{\sigma}(\vartheta_{i})&\text{ when }k\text{ is even, }n\text{ is odd}\\[15.0pt] \cfrac{1}{2^{n-1}}\sum\limits_{i=0}^{\frac{n-2}{2}}\chi_{\sigma_{i}}(\vartheta_{k})\chi_{\sigma}(\vartheta_{i})+\cfrac{1}{2^{n}}\big{(}\chi_{\sigma_{\frac{n}{2}}}(\vartheta_{k})\chi_{\sigma}(\vartheta_{\frac{n}{2}})\big{)}&\text{ when}k,n\text{ both are even}\end{cases}$
@@ 2.9. Dickson Invariants
for certain $\operatorname{GL}(E)$-invariant polynomials $d_{i}$ of degree $2^{n}-2^{n-i}$, known as Dickson invariants. Note that $w(\operatorname{reg}(E))=\mathcal{D}(E)$, where $\operatorname{reg}(E)$ is the regular representation (as mentioned in [Wil83]).
@@
so that
$\mathcal{D}(E)=\prod_{k}\mathcal{D}^{[k]}({\bm{v}}).$
Clearly, $\mathcal{D}^{[k]}({\bm{v}})$ is a symmetric polynomial. We have
$\mathcal{D}^{[1]}({\bm{v}})=1+\mathcal{E}_{1}+\mathcal{E}_{2}+\ldots+\mathcal{E}_{n},$
where $\mathcal{E}_{i}=\mathcal{E}_{i}({\bm{v}})$ is the $i$th elementary symmetric polynomial in the $v_{i}$. Also,
$\mathcal{D}^{[n]}({\bm{v}})=1+\mathcal{E}_{1}.$
If $n=4$, then
$\begin{split}\mathcal{D}^{[2]}({\bm{v}})&=1+\mathcal{E}_{1}+\mathcal{E}_{1}^{2}+\mathcal{E}_{1}^{3}+(\mathcal{E}_{2}^{2}+\mathcal{E}_{1}\mathcal{E}_{3})+(\mathcal{E}_{1}\mathcal{E}_{2}^{2}+\mathcal{E}_{1}^{2}\mathcal{E}_{3})+(\mathcal{E}_{1}\mathcal{E}_{2}\mathcal{E}_{3}+\mathcal{E}_{3}^{2}+\mathcal{E}_{1}^{2}\mathcal{E}_{4}),\\ \mathcal{D}^{[3]}({\bm{v}})&=1+\mathcal{E}_{1}+(\mathcal{E}_{1}^{2}+\mathcal{E}_{2})+(\mathcal{E}_{3}+\mathcal{E}_{1}^{3})+(\mathcal{E}_{1}^{2}\mathcal{E}_{2}+\mathcal{E}_{1}\mathcal{E}_{3}+\mathcal{E}_{4}).\end{split}$
@@
$w(\sigma)=\prod\limits_{k=1}^{n}(\mathcal{D}^{[k]}({\bm{v}}))^{m_{k}(\sigma)}$
@@ 3.1. Subgroups of $G$
meaning the nonzero entries of $A$ lie either on the diagonal or the antidiagonal. Let $Z_{X}$ be the center of $X$; it is the subgroup of diagonal matrices in $G$ which have $1$ or $-1$ on the diagonal.
@@ 3.2. Detecting SWCs
It is well known (see [FP06, Chapter VI, Sec. 5] for instance) that the mod $2$ cohomology ring of $\mathcal{S}$ is
@@
In [MS23, Corollary 4.8], we specified an orthogonal representation $\eta$ of $\mathcal{S}$ with the property that $w_{4}(\eta)=\mathfrak{e}$. We also studied the restriction map $\operatorname{H}^{*}(\mathcal{S})\to H^{*}(Z)$ and computed that $\mathfrak{e}$ maps to $v^{4}\in\operatorname{H}^{4}(Z)$. This gives an isomorphism
@@
Lemma 3.2 ([AM04], Chapter VII, Lemma 6.2).
@@
Since $\operatorname{H}^{5}(\mathcal{S})=\operatorname{H}^{6}(\mathcal{S})=\{0\}$, we necessarily have $\operatorname{Sq}^{1}(\mathfrak{e})=\operatorname{Sq}^{2}(\mathfrak{e})=0$. By naturality of $\operatorname{Sq}$ we also have $\operatorname{Sq}^{1}(\mathfrak{e}_{i})=\operatorname{Sq}^{2}(\mathfrak{e}_{i})=0$ for $i=1,2$. By Cartan’s formula (3), we deduce $\operatorname{Sq}^{2}(\mathfrak{e}_{1}^{s}\mathfrak{e}_{2}^{t})=0$ for all nonnegative $s,t$. The conclusion follows.
@@
For $k\equiv 2$ (mod $4$), we use Wu’s formula (Proposition 2.1) with $i=2$ and $j=k-2$, giving
$\begin{split}\operatorname{Sq}^{2}(w_{k-2}(\pi))&\small={k-5\choose 0}w_{2}(\pi)w_{k-2}(\pi)+{k-4\choose 1}w_{1}(\pi)w_{k-1}(\pi)+{k-3\choose 2}w_{0}(\pi)w_{k}(\pi)\\ &=\frac{(k-3)(k-4)}{2}w_{k}(\pi)\end{split}$
where we have second equality because $\operatorname{H}^{i}(\mathcal{S}\times\mathcal{S})=\{0\}$ for $i=1,2$. Also $\frac{(k-3)(k-4)}{2}$ is odd for $k\equiv 2$ (mod $4$), and $\operatorname{Sq}^{2}(w_{k-2}(\pi))=0$ by Lemma 3.3. Hence $w_{k}(\pi)=0$.
A similar argument works for $k\equiv 3$ (mod $4$): applying Wu’s formula with $i=1$ and $j=k-1$ shows that $w_{k}(\pi)=0$. ∎
We can generalize the above result to:
@@
is an orthogonal representation of $\mathcal{S}\times\mathcal{S}$. By Theorem 3.1, Proposition 3.4 and Corollary 6.1.1, we deduce $w(\Pi)\in\mathbb{F}_{2}[\mathfrak{e}_{1},\ldots,\mathfrak{e}_{n}]$.
Next we treat symmetrizations of decomposable representations. For representations $\pi_{1},\ldots,\pi_{n}$ of $\mathcal{S}$, put
$\Pi=\pi_{1}\boxtimes\cdots\boxtimes\pi_{n}.$
Then, $w(S(\Pi))=\kappa(c(\Pi))$. From Corollary 6.1.1, $c(\Pi)$ lies in the subalgebra of $\operatorname{H}^{*}(\mathcal{S}^{n},\mathbb{Z})$ generated by the $p_{j}^{*}c_{i}(\pi_{j})$. Apply the homomorphism $\kappa$; then Theorem 3.1 gives
@@
The image lands in the $S_{n}$-invariant subalgebra, as in Section 2.2. ∎

### v2

@@ Abstract.
Stiefel-Whitney classes, Symplectic Groups, Finite groups of Lie type, Weil Representations
@@ 1. Introduction
This paper is part of a project to understand SWCs for finite groups of Lie type. Let $q$ be an odd prime power throughout. A formula determining SWCs for $G=\operatorname{GL}(n,q)$ was discovered in [GJgln], for $G=\operatorname{SL}(2,q)$ in [NSSL2] and for $G=\operatorname{SL}(2n+1,q)$ in [MSSLn]. In this paper we find similar expressions when $G=\operatorname{Sp}(2n,q)$.
Write $\operatorname{H}^{*}(G)$ for the mod $2$ cohomology $\operatorname{H}^{*}(G,\mathbb{F}_{2})$, and $\operatorname{H}^{*}_{\operatorname{SW}}(G)$ for the subalgebra of $\operatorname{H}^{*}(G)$ generated by SWCs of orthogonal representations. The diagonal matrices in $G$ with eigenvalues $\pm 1$ form a subgroup we denote by $Z_{X}$. (It is the center of another subgroup $X$ which we will encounter later.) The mod $2$ cohomology of $Z_{X}$ is a polynomial algebra:
@@
Since no information is lost by restriction, we will express SWCs in terms of the $\mathcal{E}_{k}$. For $0\leq i\leq n$, let $g_{i}\in Z_{X}$ with 1-1 having multiplicity $2i$ as an eigenvalue, and $1$ having multiplicity 2(ni)2(n-i). For example $g_{0}=\mathbbm{1}$, the identity matrix. Given a representation $\pi$, write $\chi_{\pi}$ for its character. Here is our “universal” formula for the 4th and 8th SWCs:
@@
$w_{8}(\pi)=r_{1}\mathcal{E}_{2}+\left(\binom{r_{1}}{2}+\binom{r_{2}}{2}\right)\mathcal{E}_{1}^{2}\quad\text{ for $n\geq 2$}.\\$
@@
A general formula for the total SWCs $w(\pi)$ is found in Theorem 3.6, although it takes some computation to extract individual SWCs $w_{k}(\pi)$ from this. For example:
@@
Let $m\geq n$. The restriction map $\iota_{n}^{*}:\operatorname{H}^{i}(G_{m})\to\operatorname{H}^{i}(G_{n-1})$ is injective for i<4n1i<4n-1.
(Compare [NSn, Corollary 6.7] and [GJgln, Theorems 3, 8].)
@@ 2. Notations and Preliminaries
As this paper is a continuation of [NSSL2], we use the same notations and conventions, which we now review.
@@ 2.1. Representations
Let $G$ be a finite group. All the representations $(\pi,V)$ considered in this paper (before the Appendix) are complex finite dimensional. Let $\operatorname{Irr}(G)$ be the set of isomorphism classes of irreducible representations of $G$. Write $(\pi^{\vee},V^{\vee})$ for the dual representation. If $H$ is a subgroup of $G$, write $\operatorname{res}^{G}_{H}\pi$ or $\pi|_{H}$ for the restriction of $\pi$ to $H$. A linear character $\chi$ of $G$ is a degree 1 representation. We call $\chi$ quadratic when $\chi^{2}=1$.
We say $\pi$ is orthogonal (resp., symplectic), provided there exists a non-degenerate $G$-invariant symmetric (resp., antisymmetric) bilinear form $B:V\times V\to\mathbb{C}$. When $\pi$ is self-dual and irreducible, it is either orthogonal or symplectic. In this case, the Frobenius-Schur Indicator $\varepsilon(\pi)$ is a sign defined as $1$ when $\pi$ is orthogonal, and 1-1 when $\pi$ is symplectic. Whereas it is 0, when $\pi$ is not self-dual.
One can symmetrize a general $(\pi,V)$ by defining $S(\pi):=\pi\oplus\pi^{\vee}$ on the vector space $V\oplus V^{\vee}$. Under the symmetric $G$-invariant bilinear map $B$ on $V\oplus V^{\vee}$ as $B((v,\alpha),(w,\beta))=\langle\alpha,w\rangle+\langle\beta,v\rangle$, the representation $S(\pi)$ is orthogonal. We call $S(\pi)$ the symmetrization of $\pi$. (In [shl, Definition 4.3 ] it is called the hyperbolic space on $V$.)
@@ 2.3. Characteristic Classes
known as the $i$th Stiefel-Whitney Class (SWC) of $\pi$. Their sum $w(\pi)=w_{0}(\pi)+w_{1}(\pi)+\ldots$ is called the total SWC of $\pi$. We refer the reader to [NSSL2, Section 2.3] for detailed description.
Also, associated to a complex representation $\pi$ of $G$ are cohomology classes $c_{i}(\pi)\in\operatorname{H}^{2i}(G,\mathbb{Z}),$ called Chern classes (CCs). Their sum
$c(\pi)=c_{0}(\pi)+c_{1}(\pi)+c_{2}(\pi)+\ldots\in\operatorname{H}^{*}(G,\mathbb{Z})$
is called the total Chern class of $\pi$. We have
@@
where $\kappa:\operatorname{H}^{*}(G,\mathbb{Z})\rightarrow\operatorname{H}^{*}(G,\mathbb{F}_{2})$ is the coefficient homomorphism of cohomology. (See [ganguly2022stiefel, Lemma 1], based on [milnor, Problem 14-B] for proof.) So we interpret $w(S(\pi))$ as the “mod $2$ Chern class” of $\pi$.
@@
Proposition 2.1 ([may], Chapter 23, Section 6).
@@ 2.4. Symmetric Functions
Whereas for $k>n$, these are defined to be 0. For example, $\mathcal{E}_{1}(x_{1},x_{2})=x_{1}+x_{2}$ and $\mathcal{E}_{2}(x_{1},x_{2})=x_{1}x_{2}$. Of course,
@@ 2.5. Cyclic Groups
It is known [KT] that for $k\equiv 2\pmod{4}$, $\operatorname{H}^{*}(C_{k})=\mathbb{F}_{2}[v]$ where $v=w_{1}(\operatorname{sgn})$. In this case, (see [NSSL2, Lemma 2.5] for instance) for a representation $\pi$,
@@ 2.6. OIRs of a Direct Product of Groups
for $(g_{1},\ldots,g_{n})\in G$ and $v_{1}\otimes\cdots\otimes v_{n}\in V_{1}\otimes\cdots\otimes V_{n}.$ (See [fulton, Exercise 2.36, page 24].) If each $\pi_{i}$ is irreducible, then so is $\pi$. All irreducible representations of $G$ decompose in this manner. The Frobenius-Schur indicator of $\pi$ is given by
@@ 2.7. Quaternion Group
Let $Z$ be the center of $Q$, which is $\{\pm 1\}$. For a linear character $\theta$ of $Z^{n}$, write $\theta,\pi]$ for the multiplicity of $\theta$ in $\pi|_{Z^{n}}$.
@@
Let $\pi$ be an orthogonal representation of $Q^{n}$. Let $\theta$ be a non-trivial linear character of $Z^{n}$. Then $\theta,\pi]$ is a multiple of $4$.
@@
[…]
Then $\pi|_{Z^{n}}\cong 2^{r}\theta_{\pi}$, since $\rho|_{Z}=\operatorname{sgn}\oplus\operatorname{sgn}$. If $\pi$ is irreducible orthogonal, then $r$ is even by Lemma 2.2 and therefore $4$ divides $\theta_{\pi},\pi]$, and other $\theta,\pi]=0$.
@@ 2.8. $S_{n}$-invariant Representations of Elementary Abelian $2$-groups
is $S_{n}$-invariant, of degree $\binom{n}{k}$. For a polynomial $f$, write $f]_{i}$ for the coefficient of the degree $i$ term of $f$; in other words so that $f(x)=\sum_{i}[f]_{i}x^{i}$.
Lemma 2.4 ([GJgln], Proposition 2).
@@
where $m_{k}(\sigma)$ are certain non-negative integers. The coefficients $m_{k}(\sigma)$ can be expressed in terms of character values of $\sigma$ at the elements of $E$.
For $0\leq k\leq n$, put $\vartheta_{k}=e_{1}+\ldots+e_{k}$. (In particular, $\vartheta_{0}=0$.) Then,
@@
From Lemma 2.4, the character value $\chi_{\sigma_{i}}(\vartheta_{k})$ is the coefficient of $x^{i}$ in the expression $(1-x)^{k}(1+x)^{n-k}$. Following [GJgln, Theorem 7] we have
@@
[…]
@@ 2.9. Dickson Invariants
for certain $\operatorname{GL}(E)$-invariant polynomials $d_{i}$ of degree $2^{n}-2^{n-i}$, known as Dickson invariants. Note that $w(\operatorname{reg}(E))=\mathcal{D}(E)$, where $\operatorname{reg}(E)$ is the regular representation (as mentioned in [wilk]).
@@
so that $\mathcal{D}(E)=\prod_{k}\mathcal{D}^{[k]}({\bm{v}})$. Clearly, $\mathcal{D}^{[k]}({\bm{v}})$ is a symmetric polynomial. We have
$\mathcal{D}^{[n]}({\bm{v}})=1+\mathcal{E}_{1}\quad\quad\text{ and }\quad\quad\mathcal{D}^{[1]}({\bm{v}})=1+\mathcal{E}_{1}+\mathcal{E}_{2}+\ldots+\mathcal{E}_{n},$
For $n=4$, we have
$\begin{split}\mathcal{D}^{[2]}({\bm{v}})&=1+\mathcal{E}_{1}+\mathcal{E}_{1}^{2}+\mathcal{E}_{1}^{3}+(\mathcal{E}_{2}^{2}+\mathcal{E}_{1}\mathcal{E}_{3})+(\mathcal{E}_{1}\mathcal{E}_{2}^{2}+\mathcal{E}_{1}^{2}\mathcal{E}_{3})+(\mathcal{E}_{1}\mathcal{E}_{2}\mathcal{E}_{3}+\mathcal{E}_{3}^{2}+\mathcal{E}_{1}^{2}\mathcal{E}_{4})\text{ and }\\ \mathcal{D}^{[3]}({\bm{v}})&=1+\mathcal{E}_{1}+(\mathcal{E}_{1}^{2}+\mathcal{E}_{2})+(\mathcal{E}_{3}+\mathcal{E}_{1}^{3})+(\mathcal{E}_{1}^{2}\mathcal{E}_{2}+\mathcal{E}_{1}\mathcal{E}_{3}+\mathcal{E}_{4}).\end{split}$
@@
$w(\sigma)=\prod\limits_{k=1}^{n}\mathcal{D}^{[k]}({\bm{v}})^{m_{k}(\sigma)}$
@@ 3.1. Subgroups of $G$
meaning the nonzero entries of $A$ lie either on the diagonal or the antidiagonal. Let $Z_{X}$ be the center of $X$; it is the subgroup of diagonal matrices in $G$ which have $1$ or 1-1 on the diagonal.
@@ 3.2. Detecting SWCs
It is well known (see [FP, Chapter VI, Sec. 5] for instance) that the mod $2$ cohomology ring of $\mathcal{S}$ is
@@
In [NSSL2, Corollary 4.8], we specified an orthogonal representation $\eta$ of $\mathcal{S}$ with the property that $w_{4}(\eta)=\mathfrak{e}$. We also studied the restriction map $\operatorname{H}^{*}(\mathcal{S})\to H^{*}(Z)$ and computed that $\mathfrak{e}$ maps to $v^{4}\in\operatorname{H}^{4}(Z)$. This gives an isomorphism
@@
Lemma 3.2 ([Milgram], Chapter VII, Lemma 6.2).
@@
Since $\operatorname{H}^{5}(\mathcal{S})=\operatorname{H}^{6}(\mathcal{S})=\{0\}$, we necessarily have $\operatorname{Sq}^{1}(\mathfrak{e})=\operatorname{Sq}^{2}(\mathfrak{e})=0$. By naturality of $\operatorname{Sq}$ we also have $\operatorname{Sq}^{1}(\mathfrak{e}_{i})=\operatorname{Sq}^{2}(\mathfrak{e}_{i})=0$ for $i=1,2$. By Cartan’s formula (3), we deduce $\operatorname{Sq}^{2}(\mathfrak{e}_{1}^{s}\mathfrak{e}_{2}^{t})=0$ for all nonnegative $s,t$. The conclusion follows. ∎
@@
For $k\equiv 2$ (mod $4$), we use Wu’s formula (Proposition 2.1) with $i=2$ and j=k2j=k-2, giving
$\begin{split}\operatorname{Sq}^{2}(w_{k-2}(\pi))&\small={k-5\choose 0}w_{2}(\pi)w_{k-2}(\pi)+{k-4\choose 1}w_{1}(\pi)w_{k-1}(\pi)+{k-3\choose 2}w_{0}(\pi)w_{k}(\pi)\\ &=\frac{(k-3)(k-4)}{2}w_{k}(\pi).\end{split}$
The second equality holds because $\operatorname{H}^{i}(\mathcal{S}\times\mathcal{S})=\{0\}$ for $i=1,2$. Also $\frac{(k-3)(k-4)}{2}$ is odd for $k\equiv 2$ (mod $4$), and $\operatorname{Sq}^{2}(w_{k-2}(\pi))=0$ by Lemma 3.3. Hence $w_{k}(\pi)=0$.
A similar argument works for $k\equiv 3$ (mod $4$): applying Wu’s formula with $i=1$ and j=k1j=k-1 shows that $w_{k}(\pi)=0$. ∎
@@
is an orthogonal representation of $\mathcal{S}\times\mathcal{S}$. By Theorem 3.1, Proposition 3.4 and Corollary A.1.1, we deduce $w(\Pi)\in\mathbb{F}_{2}[\mathfrak{e}_{1},\ldots,\mathfrak{e}_{n}]$.
Next we treat symmetrizations of decomposable representations. For representations $\pi_{1},\ldots,\pi_{n}$ of $\mathcal{S}$, put $\Pi=\pi_{1}\boxtimes\cdots\boxtimes\pi_{n}$. Then, $w(S(\Pi))=\kappa(c(\Pi))$. From Corollary A.1.1, $c(\Pi)$ lies in the subalgebra of $\operatorname{H}^{*}(\mathcal{S}^{n},\mathbb{Z})$ generated by the $p_{j}^{*}c_{i}(\pi_{j})$. Apply the homomorphism $\kappa$; then Theorem 3.1 gives
@@
The image lands in the $S_{n}$-invariant subalgebra, as in Section $2.$ ∎